y = 2px + py^2


1
Expert's answer
2021-06-21T11:19:07-0400

given Differential equation: y = 2px + py2

where p = "\frac{dy}{dx}"

by putting this we get, y = 2"\frac{dy}{dx}"x + "\frac{dy}{dx}"y2

"\implies" y = "\frac{dy}{dx}"(2x+y2)


"\implies" "\frac{dx}{dy}" y = (2x+y2)


dividing by y both sides, we get


"\implies" "\frac{dx}{dy}" = "\frac{2x}{y}" + y


"\implies" "\frac{dx}{dy}" "-" "\frac{2x}{y}" = y


now it is a linear differential equation,


integrating factor (I.F) = "e^{\int \frac{-2}{y} dy}" = "e^{-2\int \frac{1}{y} dy}" = "e^{-2ln(y)}" = "e^{ln(\frac{1}{y^2})}" = "\frac{1}{y^2}"


therefore, the solution of given differential equation is

"\implies" x"\cdot" "\frac{1}{y^2}" = "\int \frac{1}{y^2} \cdot y \cdot dy" + C where C is any constant


"\implies" "\frac{x}{y^2}" = "\int \frac{1}{y} \cdot dy" + C


"\implies" "\frac{x}{y^2}" = "ln(y)" + C


"\implies" x = y2ln(y) + Cy2